Two kinds of supply-demand situation are as follows:
D = 12 – 3P; S = -10 + 2P
D = 16 – 2P; S = 20 – 2P
where D, S and P stand for demand curve, supply curve and market price, respectively.

Which situation has a market equilibrium solution?
1
Expert's answer
2014-10-15T14:19:36-0400
The market equilibrium exists when
D = S
For the first situation
12 – 3P = -10 + 2P
the solution is:
5P = 22
P = 4,4
D = S = -1.2
For the second one
16 – 2P = 20 – 2P
0P = 4
there is no solution